Exercise 10-24 (Part Level Submission) Crane Company receives $385,000 when it issues a $385,000, 8%, mortgage note payable to finance the construction of a building at December 31, 2017. The terms provide for annual installment payments of $44,979 on December 31. Collapse question part (a) Prepare an amortization schedule of a mortgage note for two years. (Round answers to 0 decimal places, e.g. 5,275.) Annual Interest Period Cash Payment Interest Expense Reduction of Principal Principal Balance Issue date $enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places $enter a dollar amount rounded to 0 decimal places 12/31/18 enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places 12/31/19 enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places enter a dollar amount rounded to 0 decimal places
Answer #1
Annual Interest Period Cash Payment Interest Expense Reduction of Principal Principal Balance
Issue date 385000
12/31/18 44979 30800 14179 370821
12/31/19 44979 29666 15313 355508
Workings:
Interest Expense:
12/31/18 30800 =385000*8%
12/31/19 29666 =370821*8%